OpenStudy (anonymous):

Please help! MEDAL GIVEN! -A human embryo starts at a weight of 0.0125 milligrams. The weight increases exponentially at about 28.5% each day. a. write an exponential equation to model this situation b. Define the variables in your equation, include units c. Use your equation to predict the weight of the embryo after 7 days. Show the steps d. Use your equation and the guess and check method, predict when the embryo will be twice its initial weight. Give your answer to the nearest hour.

OpenStudy (anonymous):

HI :D

OpenStudy (anonymous):

hello haha :-)

OpenStudy (anonymous):

HI again XD

OpenStudy (luigi0210):

You have to use this if I'm not mistaken: \[\LARGE y=a(1+r)^2\] a = initial amount before measuring growth/decay r = growth/decay rate (often a percent) x = number of time intervals that have passed

OpenStudy (anonymous):

yes :)

OpenStudy (luigi0210):

Whoops, that should be an "x" not a 2

OpenStudy (anonymous):

x determines the time

OpenStudy (luigi0210):

So can you come up with the equation based on that?

OpenStudy (anonymous):

y=0.0125(1+0.285) ^x @Luigi0210

OpenStudy (luigi0210):

Right! Now you could do the rest pretty easily :)

OpenStudy (anonymous):

so we solved for a already haha

OpenStudy (anonymous):

for b how would i define the variables

OpenStudy (luigi0210):

When I gave you the standard equation, I told you what each variable meant, use that.

OpenStudy (anonymous):

oh ok! :) thanks so much for c it's asking for 7 days so i just plug 7 days into x

OpenStudy (anonymous):

i'm just not sure how to do d

OpenStudy (luigi0210):

For c, yes. And for D, take the initial weight, double it, and set it equal to the equation and solve. \[\LARGE 0.25=0.0125(1+0.285)^x\]

OpenStudy (anonymous):

oh ok :) here I believe we use logs… to estimate the value of x

OpenStudy (luigi0210):

That's actually the EXACT way to find it.. the guess and check way is to just plug in random values and see which gives you one closest to 0.25

OpenStudy (anonymous):

so here we would actually divide by 0.0125 both sides of the equal sign.. how would we plug in for logs?

OpenStudy (isaiah.feynman):

When the log part comes use this!